OpenStudy (anonymous):

Part A: Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days. ANSWER: f(x)= 85x+55 charge a fee of 55 bucks + 85 days per day of rental Part B: Write the equation obtained in Part A using function notation. ANSWER: {x + 2 y = 225, x + 5 y = 480} {x = 55, y = 85} Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (anonymous):

I Need help with part C please

OpenStudy (anonymous):

@Johnbc @hartnn

OpenStudy (triciaal):

Part A minor correction charge a fee of 55 bucks + 85 days per day of rental should be $85/day + $55 The question said write the equation in standard form, Ax + By = C therefore total cost of rental - $85/per day = $55 y - 85 x = 55

OpenStudy (triciaal):

Part B f(x) =85 x + 55 { x: x>0 } y >55 f(0)=55 f(2) = 225 f(5) = 480 (2, 225) (5, 480) what is set notation? could you explain this? {x = 55, y = 85}

OpenStudy (triciaal):

Part C slope = 85 y-intercept = 55 you have 3 points on the line graph showing total cost of rental , dependent variable, y vs number of days rented, independent variable, x  on the y-axis the units/scale could be $100 because for each day the added cost is $85 the x axis, days the unit is one